# Xinyuan Lu1    Liangming Pan 2    Yubo Ma3    Preslav Nakov4    Min-Yen Kan1

### v2

@@ Abstract
Current Large Language Models (LLMs) exhibit limited ability to understand table structures and to apply precise numerical reasoning, which is crucial for tasks such as table question answering (TQA) and table-based fact verification (TFV). To address these challenges, we introduce our Tool-Augmented Reasoning framework for Tables (TART), which integrates LLMs with specialized tools. TART contains three key components: a table formatter to ensure accurate data representation, a tool maker to develop specific computational tools, and an explanation generator to maintain explainability. We also present the ToolTab dataset, a new benchmark designed specifically for training LLMs in table–tool integration. Our experiments indicate that TART achieves substantial improvements over existing methods (e.g., Chain-of-Thought) by improving both the precision of data processing and the clarity of the reasoning process. Notably, TART paired with CodeLlama achieves 90.0% of the accuracy of the closed-sourced LLM GPT-3.5-turbo, highlighting its robustness in diverse real-world scenarios. All the code and data are available at https://github.com/XinyuanLu00/TART.
@@ 1 Introduction
Tabular data is prevalent across multiple fields such as scientific research, financial reporting, and healthcare records [11]. Manual handling of such data can be both routine and error-prone, or may require specialized skills, highlighting the need for automated table reasoning [2]. Typical table-based reasoning tasks include table question answering (TQA), which extracts precise information from tables to answer given queries [7, 48, 25], and table-based fact verification (TFV), which verifies the correctness of statements by cross-referencing them with facts stored in tables [33, 22].
Figure 1: Example of the TableQA task, demonstrating the verification of travel time via boat schedule and demonstrating the critical skills needed for accurate table reasoning: table structure understanding, precise numerical calculations, and executing sequential reasoning steps.
@@
Modern large language models (LLMs) such as GPT-4 [26] have shown remarkable reasoning capabilities across a variety of tasks, spurring interest in their application to table-based tasks [41]. However, table-based reasoning presents unique challenges for LLMs, which are primarily trained on sequential text data [44], as illustrated by a real-world example in Figure 1. (1) Understanding and operating on the structure of teh table: LLMs must adapt to the non-linear format of the tables, which demands unique reasoning skills such as recognizing headers, interpreting the roles of the rows and the columns, and precisely extracting information from relevant table cells. (2) Precise numerical reasoning: Tables often contain quantitative information that requires precise calculations and comparisons. LLMs must perform operations such as summation, averaging, or trend analysis accurately, often over multiple cells or tables, which is a shift from their usual text-based reasoning tasks [15, 18]. (3) Reasoning planning: LLMs often need to plan several reasoning steps ahead. This includes decomposing the original question, determining which table parts are relevant and anticipating intermediate calculations or data transformations.
@@
Existing approaches that use LLMs for table-based reasoning can be broadly classified into two categories. One is chain-of-thought (CoT) reasoning [37], in which the model is prompted to perform step-by-step reasoning over the input table flattened as a textual sequence [44, 16, 4, 41]. Despite its flexibility, CoT often lacks precision in tabular operations and numerical reasoning, such as sorting, counting, and filtering [39]. On the other hand, program-based reasoning (PoT) [12, 5] prompts the model to generate SQL or Python code to enable precise reasoning [19, 45, 36, 39]. However, this method tends to struggle with planning and its reasoning is less understandable for humans [45]. Therefore, there is potential value in integrating the advantages of program-based and textual reasoning, to achieve both high precision and explainability in table-based reasoning.
@@
Inspired by the recent paradigm of tool-augmented language models [35, 31], we propose Tool-Augmented Reasoning framework for Tables (TART), which integrates external tool calling into the chain-of-thought reasoning process, as shown in Figure 2. Initially, TART processes the input table using a specialized module table formatter to clean and to format the raw table data, preparing it for the subsequent table operations. Subsequently, the tool maker calls specialized tools (Python functions) for tabular manipulation and numerical reasoning (e.g., adding columns, selecting rows, and grouping). Alongside these tools, TART also crafts a reasoning plan that outlines the programmatic calling sequence of the tools, specifying the necessary arguments and the expected return values for each call. Finally, following the structured reasoning plan, the explanation generator produces a hybrid text-and-program output, integrating calls to external tools into coherent, human-readable chain-of-thought explanations. Hence, TART efficiently delegates table operations and precise numerical calculations to generated tools, while preserving the planning and explainability of CoT.
@@
To train the modules in TART, we further synthesize the ToolTab dataset by distilling knowledge from a teacher LLM. We evaluate TART on nine different table-based reasoning benchmarks. The results highlight the effectiveness of integrating task-specific tools for enhancing complex reasoning capabilities. Notably, TART consistently outperformed the CoT baseline, achieving near-parity with GPT-3.5-turbo on several benchmarks, showcasing its practical utility in real-world scenarios.
@@ 2 Related Work
Table-based reasoning tasks involve interpreting and manipulating data from structured tabular sources to answer questions, verify facts, or generate summaries. Early approaches used executable SQL or SPARQL to interact with tabular data [42, 43], or graph neural networks to better encode table structures [46, 40]. However, they typically suffer from poor generalization capabilities due to their reliance on specific table formats and linguistic patterns.
Recently, large language models (LLMs) have shown great potential in improving table-based reasoning tasks. Pioneering studies have focused on developing table pre-training strategies, where LLMs are trained with sentence-table pairs to enhance their general table reasoning capabilities [6, 15, 47, 13, 41]. Subsequent work further explored different reasoning strategies to improve efficacy. For example, Ye et al.[41] proposed using LLMs as decomposers to effectively reason over tables by breaking down large tables into smaller sub-evidence and complex questions into simpler sub-questions. Want et al.[36] introduced the Chain-of-Table framework, where tabular data is explicitly used in the reasoning chain for intermediate thoughts. ReAcTable [45] combined reactive and proactive reasoning strategies to improve accuracy and retrieval from complex tables. Despite these advances, these approaches primarily rely on pure textual reasoning, such as chain-of-thought, which sacrifices the precision required for table manipulations and numerical reasoning, both crucial for table-based reasoning.
To address these limitations, TART extends the use of LLMs with integrated external tools, aimed at improving reasoning precision with function execution while maintaining explainability with LLM-based reasoning. Unlike previous methods that focus on a single reasoning strategy, TART enhances table reasoning with multiple strategies: formatting tables for better data representation, calling tools for precise calculation, and generating user-friendly explanations that clarify the reasoning process.
@@ 3 Methodology
Generally, a table-based reasoning model, $f_{\theta}(\cdot)$, parameterized by $\theta$, takes an input query $Q$ and a table $\mathcal{T}$ to produce a response $Y=f_{\theta}(Q,\mathcal{T})$. Based on this generic formulation, the nature of $Q$ and $Y$ differs depending on the specific table reasoning task: in table-based QA, $Q$ is a question and $Y$ is the answer; in table-based fact verification, $Q$ is a claim and $Y$ is its veracity label; in table summarization, $Q$ specifies the requirement and $Y$ is the summary. A table $\mathcal{T}$ is characterized by a caption $P$ and its contents ${T_{i,j}\mid i\leq R_{T},j\leq C_{T}}$, where $R_{T}$ and $C_{T}$ represent the number of rows and columns, respectively. Each cell $(i,j)$ contains data $T_{i,j}$.
@@
1. Table Formatter. TART first transforms the original table $\mathcal{T}$ with guidance from the query $Q$ into a formatted table $\mathcal{T}^{\prime}$. The formatter optimizes data formats, aligns columns, and adjusts data types as needed for the query, producing a well-formatted table that is used in subsequent reasoning.
2. Tool Maker. Given $\mathcal{T}^{\prime}$ , the tool maker generates a set of candidate tools $S$ useful for solving $Q$. It also develops a reasoning plan $R$ that details the high-level reasoning, which includes the tool calling order, as well as the necessary arguments and the expected return values for the tool calls.
3. Explanation Generator. Given the reasoning plan $R$ as a programmatic guide for chain-of-thought reasoning, the explanation generator is responsible for producing a user-friendly explanation $E$ that incorporates the use of the tools. The explanation also concludes with the final answer $\mathcal{A}$, derived from the execution of the reasoning plan $R$.
@@ 3.1 Table Formatter
We first train a specialized open-sourced large language model as the table formatter $\mathcal{F}$, which transforms the noisy raw input table $\mathcal{T}$, into a more structured and manageable format, $\mathcal{T}^{\prime}$, to facilitate subsequent reasoning
$\mathcal{T}^{\prime}=\mathcal{F}(\mathcal{T},Q)$ (1)
where the output table $\mathcal{T}^{\prime}$ is formatted in three aspects. (1) Data Cleaning: the model formats the cell values, such as removing currency symbols and textual footnotes to facilitate the execution of external functions to perform table operations or numerical reasoning. (2) Data Standardization: It converts different data representations into a uniform format, e.g., transforms the data from formats like “MM/DD/YYYY” to a consistent “YYYY-MM-DD” format across the entire table. (3) Error Handling: the model is also responsible for fixing obvious errors or missing values in the table, such as automatically inferring header names for columns without the table headers. We introduce the table formatter to ensure that the data in the input table is uniform and optimized for subsequent reasoning, especially to make it more compatible with function execution. In practice, we transform the formatted table $\mathcal{T}^{\prime}$ into a Python array, facilitating easier interpretation and processing by subsequent reasoning modules.
@@ 3.2 Tool Maker
Recent studies have shown that LLMs have the capability of synthesizing relevant tools by understanding the problem context and creating solutions based on the crafted tools [31, 3, 35]. Motivated by this, we train another specialized LLM $\mathcal{M}$ as a tool maker, which takes as input the reformatted table $\mathcal{T}^{\prime}$ and the query $Q$ to generate a set of candidate tools $S$ and develops a reasoning plan $R$ that details the high-level reasoning steps.
$S,R=\mathcal{M}(\mathcal{T^{\prime}},Q).$ (2)
The tool set $S=\{s_{1},\cdots,s_{n}\}$ consists of $n$ specialized tools and each tool $s_{i}$ is a Python function that performs table operations (e.g., get_column_by_name), numerical reasoning (e.g., average, argmax), or higher-level functions (e.g., linear_regression). In table-based reasoning, these automated tools are essential to handle reasoning tasks that textual-based LLMs cannot address effectively, thereby significantly enhancing their problem-solving capacities.
Unlike previous work that manually defined a small number ($<10$) of hand-crafted tools [20, 27] or retrieved tools from a predefined set [29, 23], we choose to train a specialized tool maker model that learns to generate tools dynamically, based on the specifics of the table and the context of the problem. This approach not only preserves the model’s ability to “extract” previously encountered tools from its parametric memory, but also empowers the model to create novel tools as needed for unique problems, as shown in Section 4.4. While generating tools offers greater flexibility, it is crucial to prevent the tool maker from creating overly-specific tools (e.g., count_people_on_third_floor), as this would hinder its ability to generalize to new problems. To address this issue, we incorporate tool abstraction and tool deduplication steps when constructing synthetic data for training the module (elaborated on in Section 3.4).
In addition to generating useful tools, the model also constructs a high-level reasoning plan $R=[r_{1},\cdots,r_{N}]$, which outlines how the tools should be applied. The reasoning plan is formulated as a sequence of $N$ function calls. Each function call $r_{i}=(s_{i},A_{i},V_{i})$ includes the function $s_{i}\in\mathcal{S}$, the argument $A_{i}$ passed to the function, and the variable $V_{i}$ that stores the result of the function call $s_{i}(A_{i})$. This reasoning plan acts as a programmatic blueprint, guiding the table-based reasoning process. Both the tool set $S$ and the reasoning plan $R$ are then provided to the explanation generator, which produces the final explainable reasoning output.
@@ 3.3 Explanation Generator
While program-based reasoning plans are precise, they are often difficult for non-expert users to understand. Moreover, certain types of reasoning, such as commonsense or narrative-based reasoning, are better communicated in natural language. To address this, TART incorporates a specialized module called the explanation generator $\mathcal{E}$, which generates chain-of-thought natural language explanations integrated with function calls, following the steps outlined in the reasoning plan $R$
$O=\mathcal{E}(S,R).$ (3)
The final output $O$ of TART provides detailed explanations for the function calls. For example, the function call get_column_by_name is explained as, “First, retrieve the column listing snowfall in inches.” Additionally, the explanation generator groups related function calls together to create coherent and easy-to-follow explanations, as illustrated in Figure 2.

### v3

@@ Abstract
Current Large Language Models (LLMs) exhibit limited ability to understand table structures and to apply precise numerical reasoning, which is crucial for tasks such as table question answering and table-based fact verification. To address these challenges, we introduce our Tool-Augmented Reasoning framework for Tables (TART), which integrates LLMs with specialized tools. TART contains three key components: a table formatter to ensure accurate data representation, a tool maker to develop specific computational tools, and an explanation generator to maintain explainability. We also present the ToolTab dataset, a new benchmark designed specifically for training LLMs in table–tool integration. Our experiments indicate that TART achieves substantial improvements over existing methods (e.g., Chain-of-Thought) by improving both the precision of data processing and the clarity of the reasoning process. Notably, TART paired with CodeLlama achieves 90.0% of the accuracy of the closed-sourced LLM GPT-3.5-turbo, highlighting its robustness in diverse real-world scenarios. Both code and data are openly available at https://github.com/XinyuanLu00/TART.
@@ 1 Introduction
Tabular data is prevalent across multiple fields such as scientific research, financial reporting, and healthcare records Dong et al. (2022). Manual handling of such data can be both routine and error-prone, or may require specialized skills, highlighting the need for automated table reasoning to improve efficiency Badaro et al. (2023). Typical table-based reasoning tasks include table question answering (TQA), which extracts precise information from tables to answer given queries Chen et al. (2020b); Zhu et al. (2021); Nan et al. (2022) and table-based fact verification (TFV), which verifies the correctness of statements by cross-referencing them with facts stored in tables Wang et al. (2021); Lu et al. (2023c).
Figure 1: Example of the TableQA task, demonstrating the verification of travel time via boat schedule and the critical skills needed for accurate table reasoning: table structure understanding, precise numerical calculations, and executing sequential reasoning steps.
@@
Modern large language models (LLMs) such as GPT-4 OpenAI (2023) have shown remarkable reasoning capabilities across a variety of tasks, spurring interest in their application to table-based tasks Ye et al. (2023). However, table-based reasoning presents unique challenges for LLMs, which are primarily trained on sequential text data Zhang et al. (2024a), as illustrated by a real-world example in Figure 1. (1) Understanding and operating on table structure: LLMs must adapt to the non-linear format of the tables, which demands unique reasoning skills such as recognizing headers, interpreting the roles of the rows and the columns, and precisely extracting information from relevant table cells. (2) Precise numerical reasoning: Tables often contain quantitative information that requires precise calculations and comparisons. LLMs must perform operations such as summation, averaging, or trend analysis accurately, often over multiple cells or tables, which is a shift from their usual text-based reasoning tasks Herzig et al. (2020); Liu et al. (2022). (3) Reasoning planning: LLMs often need to plan several reasoning steps ahead. This includes decomposing the original question, determining relevant table parts, and anticipating intermediate calculations or data transformations. All three challenges manifest in Figure 1.
@@
Existing approaches that use LLMs for table-based reasoning can be broadly classified into two categories. One is chain-of-thought (CoT) reasoning Wei et al. (2022), in which the model is prompted to perform step-by-step reasoning over the input table flattened as a textual sequence Zhang et al. (2024a); Jin and Lu (2023); Chen (2023); Ye et al. (2023). Despite its flexibility, CoT often lacks precision in tabular operations and numerical reasoning, such as sorting, counting, and filtering Wu and Feng (2024). The second approach, program-based reasoning (PoT) Gao et al. (2023); Chen et al. (2023) prompts the model to generate SQL or Python code to enable precise reasoning Liu et al. (2024); Zhang et al. (2024b); Wang et al. (2024b); Wu and Feng (2024). However, this method struggles with reasoning planning and its reasoning is less understandable to humans Zhang et al. (2024b). Therefore, there is potential value in integrating the advantages of program-based and textual reasoning, to achieve both high precision and explainability in table-based reasoning.
@@
Inspired by the recent paradigm of tool-augmented language models Wang et al. (2024a); Schick et al. (2023), we propose Tool-Augmented Reasoning framework for Tables (TART), which integrates external tool calling into the chain-of-thought reasoning process, as shown in Figure 2. Initially, TART processes the input table using a specialized module table formatter to clean and to format the raw table data, preparing it for the subsequent table operations. Subsequently, the tool maker calls specialized tools (Python functions) for tabular manipulation and numerical reasoning (e.g., adding columns, selecting rows, and grouping). Alongside these tools, TART also crafts a reasoning plan that outlines the programmatic calling sequence of the tools, specifying the necessary arguments and the expected return values for each call. Finally, following the structured reasoning plan, the explanation generator produces a hybrid text-and-program output, integrating calls to external tools into coherent, human-readable chain-of-thought explanations. In doing so, TART efficiently delegates table operations and precise numerical calculations to generated tools while preserving CoT’s planning ability and explainability.
@@
To train the modules in TART, we further synthesize the ToolTab dataset by distilling knowledge from a teacher LLM. We evaluate TART on nine different table-based reasoning benchmarks. The results highlight the effectiveness of integrating task-specific tools for enhancing complex reasoning capabilities. Notably, TART consistently outperformed the CoT baseline, achieving near-parity with GPT-3.5-turbo on benchmarks, showcasing its usefulness in real-world scenarios.
In summary, our contributions are threefold:
$\bullet$ We propose TART, a novel framework that enhances table-based reasoning by integrating tools into the reasoning process, which addresses the limitations of current LLMs in handling table structures and executing precise calculations.
$\bullet$ We develop ToolTab, a comprehensive benchmark specifically designed to train LLMs on table-tool integration. It includes diverse real-world tables, uniform format, and careful validation to ensure high-quality training.
$\bullet$ Our experiments confirm that TART not only improves the precision and the explainability of table-based reasoning, but also generalizes effectively to out-of-domain datasets.
@@ 2 Related Work
Table-based reasoning tasks involve interpreting and manipulating data from structured tabular sources to answer questions, verify facts, or generate summaries. Early approaches used executable SQL or SPARQL to interact with tabular data Yin et al. (2016); Yu et al. (2018), or graph neural networks to better encode table structures Zhong et al. (2020); Yang et al. (2020). However, they typically suffer from poor generalization capabilities due to their reliance on specific table formats and linguistic patterns.
Recent advances in large language models (LLMs) have demonstrated significant potential in this area. Pre-training strategies that align LLMs with sentence-table pairs Chen et al. (2020a); Herzig et al. (2020); Zhou et al. (2022); Gu et al. (2022); Ye et al. (2023); Glockner et al. (2024) have improved table reasoning capabilities, while frameworks like TAP4LLM Sui et al. (2024b) optimizes table representations through sampling, augmentation, and serialization. Other works, such as ReAcTable Zhang et al. (2024b) and Chain-of-Table Wang et al. (2024b), have introduced hybrid or explicit reasoning mechanisms to better integrate tabular data into reasoning chains. These methods primarily rely on textual reasoning strategies, such as chain-of-thought, which often lack the precision necessary for table manipulations and numerical reasoning. Efforts have also been made to evaluate and enhance LLM capabilities for table-related tasks. Studies like Table Meets LLM Sui et al. (2024a) and Text2Analysis He et al. (2024) introduce benchmarks for tasks such as cell lookup, row retrieval, and Python-based advanced data analysis, emphasizing challenges in table serialization and query understanding. Meanwhile, TabularNet Du et al. (2021) proposes novel architectures, combining graph-based and relational representations to improve semantic understanding of tabular data. However, these approaches typically focus on either enhancing representation or evaluation rather than augmenting reasoning precision. In addition to general-purpose frameworks, domain-specific solutions such as EHRAgent Shi et al. (2024) have been developed for multi-tabular reasoning in specialized domains like electronic health records (EHRs). While EHRAgent integrates domain-specific metadata and debugging strategies, its focus contrasts with TART, which is designed as a domain-agnostic solution for table reasoning across diverse contexts. Similarly, API-Assisted Code Generation Cao et al. (2023) translates queries into Python programs leveraging fixed APIs, differing from TART ’s dynamic tool-augmented reasoning approach.
To address the limitations of prior work, TART extends the use of LLMs with integrated external tools, enabling precise table manipulations and numerical reasoning while maintaining explainability. By combining strategies such as table formatting for better representation, tool-based function execution for precision, and LLM-based reasoning for interpretability, TART advances the state of table-based reasoning tasks.
@@ 3 Methodology
Generally, a table-based reasoning model, $f_{\theta}(\cdot)$, parameterized by $\theta$, takes an input query $Q$ and a table $\mathcal{T}$ to produce a response $Y=f_{\theta}(Q,\mathcal{T})$. Based on this generic formulation, the nature of $Q$ and $Y$ differs depending on the specific table reasoning task:
in table-based QA, $Q$ is a question and $Y$ is the answer; in table-based fact verification, $Q$ is a claim and $Y$ is its veracity label. A table $\mathcal{T}$ is characterized by a caption $P$ and its contents ${T_{i,j}\mid i\leq R_{T},j\leq C_{T}}$, where $R_{T}$ and $C_{T}$ represent the numbers of rows and columns, respectively. Each cell $(i,j)$ contains data $T_{i,j}$.
@@
1. Table Formatter.
TART first transforms the original table $\mathcal{T}$ with guidance from the query $Q$ into a formatted table $\mathcal{T}^{\prime}$. The formatter optimizes data formats, aligns columns, and adjusts data types as needed for the query, producing a well-formatted table that is used in subsequent reasoning.
2. Tool Maker.
Given $\mathcal{T}^{\prime}$ , the tool maker generates a set of candidate tools $S$ useful for solving $Q$. It also develops a reasoning plan $R$ that details the high-level reasoning which includes the tool calling order, as well as the necessary arguments and the expected return values for the tool calls.
3. Explanation Generator.
Given the reasoning plan $R$ as a programmatic guide for chain-of-thought reasoning, the explanation generator is responsible for producing a user-friendly explanation $E$ that incorporates the use of the tools. The explanation also concludes with the final answer $\mathcal{A}$, derived from the reasoning plan $R$’s execution.
@@ 3.1 Table Formatter
We first train a specialized open-sourced large language model as the table formatter $\mathcal{F}$, which transforms the noisy raw input table $\mathcal{T}$, into a more structured and manageable format, $\mathcal{T}^{\prime}$, to facilitate subsequent reasoning: $\mathcal{T}^{\prime}=\mathcal{F}(\mathcal{T},Q)$ where the output table $\mathcal{T}^{\prime}$ is formatted according to three aspects. 1) Data Cleaning: the model formats the cell values, such as removing currency symbols and textual footnotes to facilitate the execution of external functions to perform table operations or numerical reasoning. 2) Data Standardization converts different data representations into a uniform format;e.g., transforming the data from “MM/DD/YYYY” to a consistent “YYYY-MM-DD” format across the entire table. 3) Error Handling: the model also fixes obvious errors or missing values in the table, such as automatically inferring header names for columns without the table headers.
We introduce the table formatter to ensure that the data in the input table is uniform and optimized for subsequent reasoning, especially to make it more compatible with function execution. In practice, we transform the formatted table $\mathcal{T}^{\prime}$ into a Python array, facilitating easier interpretation and processing by subsequent reasoning modules.
@@ 3.2 Tool Maker
Recent studies have shown that LLMs have the capability of synthesizing relevant tools by understanding the problem context and creating solutions based on the crafted tools Schick et al. (2023); Cai et al. (2024); Wang et al. (2024a). Motivated by this, we train another specialized LLM $\mathcal{M}$ as a tool maker, which takes as input the reformatted table $\mathcal{T}^{\prime}$ and the query $Q$ to generate a set of candidate tools $S$ and develops a reasoning plan $R$ that details the high-level reasoning steps: $S,R=\mathcal{M}(\mathcal{T^{\prime}},Q).$ The tool set $S=\{s_{1},\cdots,s_{n}\}$ consists of $n$ specialized tools, where each tool $s_{i}$ is a Python function that performs table operations (e.g., get_column_by_name), numerical reasoning (e.g., average, argmax), or higher-level functions (e.g., linear_regression). These automated tools are essential to handle reasoning tasks that textual-based LLMs cannot address effectively.
Unlike previous work that manually defined a small number ($<10$) of hand-crafted tools Lu et al. (2023a); Pan et al. (2023) or retrieved tools from a predefined set Qin et al. (2024); Ma et al. (2024), we choose to train a specialized tool maker model that learns to generate tools dynamically, based on the specifics of the table and the context of the problem. This approach not only preserves the model’s ability to “extract” previously encountered tools from its parametric memory, but also empowers the model to create novel tools as needed for unique problems, as shown in Section 5.2. While generating tools offer greater flexibility, it is crucial to prevent the tool maker from creating overly-specific tools (e.g., count_people_on_third_floor), as this would hinder its ability to generalize to new problems. To address this issue, we incorporate tool abstraction and tool deduplication steps when constructing synthetic data for training the module (Section 3.4).
The model also constructs a high-level reasoning plan $R=[r_{1},\cdots,r_{N}]$, which outlines how tools should be applied. The reasoning plan is formulated as a sequence of $N$ function calls. Each function call $r_{i}=(s_{i},A_{i},V_{i})$ includes the function $s_{i}\in\mathcal{S}$, the argument $A_{i}$ passed to the function, and the variable $V_{i}$ that stores the result of the function call $s_{i}(A_{i})$.
This reasoning plan acts as a programmatic blueprint, guiding the table-based reasoning process. Both the tool set $S$ and the reasoning plan $R$ are then provided to the explanation generator, producing the final explainable reasoning output.
@@ 3.3 Explanation Generator
While program-based reasoning plans are precise, they are often difficult for non-expert users to understand. Moreover, certain types of reasoning, such as commonsense or narrative-based reasoning, are better communicated in natural language. To address this, TART incorporates a specialized module called the explanation generator $\mathcal{E}$, which generates chain-of-thought natural language explanations integrated with function calls, following the steps outlined in the reasoning plan $R$: $O=\mathcal{E}(S,R).$ The final output $O$ of TART provides detailed explanations for the function calls. For example, the function call get_column_by_name is explained as, “First, retrieve the column listing snowfall in inches.” Additionally, the explanation generator groups related function calls together to create coherent and easy-to-follow explanations, as illustrated in Figure 2.